OpenStudy (anonymous):

Jesse kicks a soccer ball off the ground and in the air with an initial velocity of 29 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches?

OpenStudy (anonymous):

the maximum height is reached at the vertex of this parabola. the vertex is obtained from x=-b/(2a)=-29/(2*(-16))=29/32. find H(29/32) to get your maximum.

OpenStudy (anonymous):

i still dont get it im slow

OpenStudy (anonymous):

you have a calculator?

OpenStudy (anonymous):

H(29/32) = -16(29/32)^2 + 29(29/32)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

use the calculator on your computer.

OpenStudy (anonymous):

would it be 14.2